OpenStudy (poopsiedoodle):

Can someone help explain how to do quadratic equations with the quadratic formula? I'll give you a problem. See comments.

OpenStudy (poopsiedoodle):

\[\huge 0 = -4(-4t^2 -8t -1)\]

OpenStudy (hba):

Well the first step is to compare your equation with the standard form and get a,b and c \[\huge\ ax^2+bx+c=0\]

OpenStudy (poopsiedoodle):

So, how would I do that with this one? Multiply the parentheses stuff by -4?

OpenStudy (hba):

Seeing your question i would suggest you to first divide your whole equation by -4 and then compare

OpenStudy (poopsiedoodle):

\[\Huge 0=-4t^2 -8t-1\]

OpenStudy (hba):

What is a,b and c in this ?

OpenStudy (poopsiedoodle):

a=-4 b=-8 c=-1

OpenStudy (hba):

Good,Now you can use the quad formula. \[\huge\ x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (hba):

The other think which you can do with your equation is that you can take -1 common and then compare so everything becomes positive.you can also do it straight away :)

OpenStudy (poopsiedoodle):

so, t = -8*√64-4*-4 ------------------ 2*-4 ?

OpenStudy (poopsiedoodle):

which would then be t=-8*(8+4) ----------- -8

OpenStudy (poopsiedoodle):

am I getting it right?

OpenStudy (hba):

It should be +- not *

OpenStudy (poopsiedoodle):

I don't get it... wouldn't you be multiplying -8 and \[\pm √64-4*-4\]?

OpenStudy (hba):

t=-8(+-)(8+4) ----------- -8

OpenStudy (poopsiedoodle):

Now I'm rather confused.

OpenStudy (hba):

Lemme show you.

OpenStudy (hba):

I appreciate you working with me though.

OpenStudy (poopsiedoodle):

I guess I could do that \(\Huge\mathsf{:P}\)

OpenStudy (hba):

a=-4 b=-8 c=-1 \[\huge\ t=\frac{ -(-8) \pm \sqrt{(-8)^2-4(-4)(-1)} }{ 2(-4) }\]

OpenStudy (poopsiedoodle):

Ok, but what does the \[\pm\] do? what's its significance?

OpenStudy (hba):

Well when we simplify it we would be taking the adding part on one side and subracting part on the other side.

OpenStudy (poopsiedoodle):

so it'd be 8+√blahblahblah --------------- 2-(-4)?

OpenStudy (hba):

Either, 8+√blahblahblah --------------- 2-(-4)? Or, 8-√blahblahblah --------------- 2-(-4)?

OpenStudy (poopsiedoodle):

so the \[\pm\] just breaks up between the denominator and the numerator?

OpenStudy (hba):

Exactly.

OpenStudy (poopsiedoodle):

Ah. I guess I'll try another problem here and ask you if it's right or something.

OpenStudy (hba):

Sure :)

OpenStudy (poopsiedoodle):

\[\huge 0=-2x ^{2} +14x +20\] \[\large x=\frac{ -14\pm \sqrt{14^2 - 4(-2+20)} }{ -4 }\] \[\large x= \frac{ -14+14-72 }{ -4 }\]

OpenStudy (poopsiedoodle):

and the answer is \[\huge \frac{ -7\pm \sqrt{89} }{ 2 }\]. Would I be right?

OpenStudy (hba):

Its is -4(a)(c) not -4(a+c)

OpenStudy (poopsiedoodle):

oh. Well that ruins that then.

OpenStudy (hba):

Yes it does :) Do it on a paper and tell me what you got from that equation :)

OpenStudy (poopsiedoodle):

\[\Large x=\frac{ -14\pm \sqrt{14^2 - 4(-2*20)} }{ -4 }\] \[\Large x=\frac{ -14\pm 14 - 4(-40)} { -4 }\] \[\Large x=\frac{ -14\pm 14-160} { -4}\] \[\Large x=\frac{ -14\pm -146} {-4}\] set x to 0 and divide all by -2 to get \[\Large \frac{ 7\pm 73} { 2}\]

OpenStudy (poopsiedoodle):

and now I'm confused.

OpenStudy (poopsiedoodle):

oh wait, I forgot to keep the x's in there. Crap.

OpenStudy (poopsiedoodle):

Think you could help me out here? My mind is about to implode.

OpenStudy (hba):

x=7+73/2,x=7-73/2

OpenStudy (hba):

Btw, (-) *(-)=+

OpenStudy (hba):

So just because of signs it goes wrong.

OpenStudy (hba):

\[\Large x=\frac{ -14\pm 14 - 4(-40)} { -4 }\] Continue from here.

OpenStudy (poopsiedoodle):

But I left the xs out of my work. ._.

OpenStudy (hba):

No worries about that :)

OpenStudy (hba):

Continue working from here.

OpenStudy (poopsiedoodle):

\[\Large x=\frac{ -14\pm 14 - 4(-40)} { -4 }\] \[\Large x=\frac{ -14\pm 14+160} {-4}\] \[\Large x=\frac{ -14\pm 174} {-4}\] \[\Large x=\frac {\frac{ -14\pm 174} {-4}} {-2}\] \[\Large x=\frac {7\pm -87} {2}\]

OpenStudy (poopsiedoodle):

And now what do I do?

OpenStudy (hba):

I don't know what you are upto ,where did that divide by -2 come in between ?

OpenStudy (poopsiedoodle):

Well, I'm not sure. I just thought I had to do that. I'm rather confused ._.

OpenStudy (hba):

Hmm,If OS would be working on chrome i would have worked one problem for you. Can we do this later ?

OpenStudy (poopsiedoodle):

I'm using chrome right now o_o But, I guess that'd be OK :P

OpenStudy (hba):

Thanks.

OpenStudy (poopsiedoodle):

Thanks to you too :3